To measure the gravity on a new planet, scientist constructed a tower and timed how long an object took to hit the surface of the planet after being dropped. If the tower was 50 meters high and the object took 4 seconds to fall, what is the acceleration due to gravity of the plane
1
Expert's answer
2020-12-14T07:30:28-0500

The acceleration due to gravity of the planet is


"g=\sqrt{2h/t^2}=\sqrt{2\cdot50/4^2}=2.5\text{ m/s}^2."